OpenStudy (anonymous):

Describe two different methods to estimate 1.35 + 2.4 + 1.48, and give their results. How do these compare to the actual solution?

OpenStudy (anonymous):

@sptogl

OpenStudy (sptogl):

1.4+2.4+1.5=5.3

OpenStudy (sptogl):

1+2+3=6

OpenStudy (anonymous):

okay

OpenStudy (sptogl):

1.35+2.4+1.48=5.23

OpenStudy (anonymous):

otay

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@Irrati0nal